OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME!! Convert x 2 + y 2 = 8 to a polar equation. r = 4 r² = 8 r² = 64

OpenStudy (anonymous):

is r the given?

OpenStudy (anonymous):

yes r is 8 its based off the formula x^2+y^2=8

OpenStudy (anonymous):

i think its the last one r^2=64 r=8

OpenStudy (anonymous):

x^2+y^2=8 ==> r^2=8

OpenStudy (anonymous):

how can r^2 be both 8 and 64?

OpenStudy (anonymous):

r^2=8

OpenStudy (anonymous):